OpenStudy (anonymous):

Solve the equation of log (2x + 4) = log (3x -12) With explanation please.

OpenStudy (austinl):

\[\log(2x+4) - \log(3x-12) = 0\] \[\log(\frac{2x+4}{3x-12})=0\] \[\frac{2x+4}{3x-12} = 1\] 2x+4 = 3x-12 -x = -16 x = 16